The length of a rectangle is five meters more than twice the width. The perimeter is 70 meters. Find the width and length.
Answer by josgarithmetic(39617)

Length L, width w.

L=5+2w and 2w+2L=70.

Simplify the perimeter equation to w+L=35.
Substitute as w+5+2w=35.
Solve first for w.